Convert the following number of moles of chemical into its corresponding mass in grams.





1. 0.436 moles of ammonium chloride





2. 2.360 moles of lead (II) oxide





3. 0.031 moles of aluminium iodide





4. 1.077 moles of magnesium phosphate





5. 0.50 moles of calcium nitrate

1
Expert's answer
2022-03-10T10:44:38-0500

"Mass= moles\u00d7 rfm"

1)rfm= 53.5

Mass= 0.436× 53.5= 23.3g

2)rfm= 223

Mass= 2.36× 223= 526.28g

3) rfm= 407.7

Mass= 0.03× 407.7= 12.64g

4) rfm=263

Mass= 263× 1.077= 283.3g

5)rfm= 164

Mass= 0.5× 164= 82g